How many cups of a 10% acid solution and a 25% solution need to be mixed to make 6 cups of a 15% solution?
Answer by stanbon(75887)

How many cups of a 10% acid solution and a 25% solution need to be mixed to make 6 cups of a 15% solution?
---------------
Equation:
acid + acid = acid
0.10x + 0.25(6-x) = 0.15*6
-----
10x + 25*6 - 25x = 15*6
----
-15x = -10*6
x = (2/3)6
x = 4 cups (amt. of 10% solution needed)
6-x = 2 cups (amt. of 25% solution needed)
---------------------------------------
Cheers,
Stan H.
-------------